OpenStudy (anonymous):

Which of the following fractions can be used in the conversion of 25 m3 to the unit cm3?

OpenStudy (anonymous):

fraction cubed one m over ten to the negative second power cm fraction cubed one cm over ten to the negative second power m one m over ten to the negative second power cm one cm over ten to the negative second power m

OpenStudy (anonymous):

fraction cubed one cm over ten to the negative second power m see 1 m = 10^2 cm hence 25 m3 =25*(10^2) cm3 =25*10^6 cm3